A small positively-charged particle that has a charge of 1.20x10^-4 C came into the vicinity of a large positively-charged particle that created an electric force upon the test charge with an amount of 23.0x10 N. What is the Electric Field Strength exerted by the point charge then?


1
Expert's answer
2021-02-22T10:19:56-0500

charge - "q = 1.2*10^{-4}C" Force - "F = 230N"

The Filed Strength exerted by the point is "E = \large\frac{F}{q}= \frac{230N}{1.2*10^{-4}C}" "= 19.1(6)*10^{5}\large\frac{N}{C}"